# TransDiffuser: End-to-end Trajectory Generation with Decorrelated Multi-modal Representation for Autonomous Driving

### v1

@@ 4.4 Qualitative Analysis
We provide the visualization of representative examples from the BEV perspective. For simple traffic scenarios where there are less notable objects on the road, our framework can propose relatively more aggressive feasible planning trajectories. For complicated traffic scenarios where there are numerous notable objects around the ego driving agent, our framework tends to propose diverse yet feasible planning trajectories on the premise of ensuring safety.
Figure 3: Visualization of representative examples of Transfuser [8] and our proposed framework.
5 Conclusion
@@
In this work, we propose TransDiffuser, an encoder-decoder based generative trajectory model for end-to-end autonomous driving. The encoded scene features and motion features serves as conditional input of the Diffusion-based denoising decoder. We introduce a simple yet effective multi-modal representation decorrelation optimization mechanism to encourage more diverse trajectories from the continuous action space. Experiments on the planning-oriented NAVSIM benchmark demonstrate its superiority in generating high-quality diverse trajectories. For future works, considering more real-world autonomous driving scenarios, the autonomous vehicle should generate multiple planning trajectories better aligned with human driver commands or styles, these attempts may require reinforcement learning techniques [9] and vision-language-action model architectures [51, 52].
6 Limitations
In our experiments, we find there exists some difficulties when finetuning the full model (i.e. including the perception encoder), which requires more technical explorations to fully unleash the potential of the proposed framework. Future works can also consider applying our introduced plug-and-play decorrelation regularization on other models or try more advanced diffusion policies which may possibly improve the performance to enrich this study.
Acknowledgments
@@
This work is a follow-up work of our team’s previous proposed TrajHF [9]. We sincerely appreciate the assistance of Jianwei Ren from Shanghai Qi Zhi Institute, along with Yue Wang, Chuan Tang and other engineers from LiAuto for TrajHF [9] re-implementation and other technical supports.

### v2

@@ IV-E Qualitative Analysis
Comparison with Transfuser. Since TransDiffuser endows the auto-regressive based Transfuser [7] with generation capabilities to some extent, we provide the visualization of representative examples from the BEV perspective. We visualize the single auto-regressive trajectory predicted by Transfuser and the single selected trajectory of our model. For simple traffic scenarios where there are fewer notable objects, TransDiffuser can propose relatively more aggressive feasible planning trajectories. For complicated traffic scenarios where there are numerous notable objects around the ego driving agent, it tends to propose diverse yet feasible planning trajectories on the premise of ensuring safety.
Figure 4: Visualization of single-mode trajectories.
Comparison with GoalFlow. Since diffusion based counterparts like GoalFlow [9] do not provide specific scene tokens for visualization, we take similar scenes as examples to compare the diversity of generated trajectories in Figure 5. Though we sampled 30 trajectories (fewer than 128/256 trajectories in GoalFlow), we can get more diverse yet feasible trajectories covering more plausible driving space. In contrast, trajectories generated by GoalFlow converge to the similar trajectory distribution given the guidance of goal point priors. By exploiting decorrelation optimization to enrich latent representation space, we can explore more diverse trajectories.
Figure 5: Visualization on multi-mode diverse trajectories.
V Conclusion
@@
In this work, we propose TransDiffuser, an encoder-decoder based generative trajectory model for end-to-end autonomous driving. The encoded scene features and motion features serves as conditional input of the Diffusion-based denoising decoder. We introduce a simple yet effective multi-modal representation decorrelation optimization mechanism to encourage more diverse trajectories from the continuous action space. Experiments on the planning-oriented NAVSIM benchmark demonstrate its superiority in generating high-quality diverse trajectories, without any trajectory anchors or scene priors. For future works, considering more real-world scenarios, the vehicle should generate multiple planning trajectories better aligned with human driver commands or styles, these attempts require reinforcement learning based preference optimization techniques [8, 23] and vision-language-action model architectures [54, 55] to achieve the better balance between diversity and safety, which is a crucial concern for deployment. We also plan to explore vision-only generative approaches on NAVSIMv2 [56] for open-looped evaluation.
ACKNOWLEDGMENT
@@
This work is a follow-up work of our team’s previous proposed TrajHF [8]. We appreciate the valuable feedback and advice from Tianyu Li from OpenDrive Lab and Hangjie Mo from Hefei University of Technology. We also appreciate the assistance of Jianwei Ren from Shanghai Qi Zhi Institute, along with Yue Wang, Chuan Tang and other researchers and engineers from LiAuto for TrajHF re-implementation and other technical supports.